Flamo:

Which of the following sets could be the sides of a right triangle? {2, 3, } (Sqare root of 10) {3, 5, } (Square Root 34) {5, 8, 12}

Flamo:

Im sTuck.

Flamo:

O: is it 2,3 and Square root of 10?

Angle:

hmmm well 2^2 + 3^2 = 4 + 9 = 13 which is not 10

Flamo:

o

Flamo:

3^2 + 5^2 = 9 + 25

Flamo:

9 + 25 = 34

Flamo:

Right?

Angle:

yup :)

Flamo:

Ye!

Flamo:

Thank You!